The degeneracy of the level of hydrogen atom that has energy -RH/16 is?


(A)       16                                            


(B)       4


(C)       2                                             


(D)       1


Please respond soon.....

Grade:12

2 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

Compare it with RH/n2 we find that n2=16 hence n=4

So the correct option is b)

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

Sujit Kumar
5 years ago
Compare it with RH/(n)^2 we find that (n)^2=16 hence n=4
For n=4 the possible orbitals are s, p, d, f
no of degenerate orbitals is the total no. Of m values
no. Of m values for 
s – 1
p – 3
d – 5
f – 7
Therefore total number of m values = 1 + 3 + 5 + 7 = 16
 
Ans: (A) 16